# Tin-Based Chalcogenide Perovskites: A Promising Lead-Free Alternative for Stable and High-Performance Photovoltaics

### v1

@@ Abstract
Chalcogenide perovskites (CPs) have sparked interest as promising optoelectronic materials due to their stability, nontoxicity, small bandgaps, large absorption coefficients, and high defect tolerance. Here, using state-of-the-art first-principles-based density functional theory, density functional perturbation theory, and many-body perturbation theory (i.e., GW and BSE), we explicate the excitonic and polaronic phenomena as well as the relative stability and optoelectronic properties in a series of distorted CPs ASnX3 (A = Ca, Sr, Ba; X = S, Se). Our findings reveal that these perovskites are mechanically stable and exhibit direct G0W0 bandgaps ranging from 0.79 to 1.50 eV. Moreover, we find that the exciton binding energy of these compounds (0.040.23 eV) is comparable to that of Zr- and Hf-based CPs but little higher than that of conventional lead halide perovskites (HPs). Additionally, we look into polaron-facilitated charge carrier mobility for electrons (21.33416.02 cm2V-1s-1) and holes (7.02260.69 cm2V-1s-1), which are comparable to or higher than those observed in lead HPs and significantly exceed those in Zr- and Hf-based CPs, owing to reduced carrier-phonon couplings in the former. Finally, the estimated spectroscopic limited maximum efficiency (24.2%31.2%) reflects that they can be promising candidates for photovoltaic applications. This has been further confirmed by assessing the performance of perovskite solar cells through conventional device (FTO/TiO2/ASnX3/Spiro-OMeTAD/Au) simulations using SCAPS-1D software.
@@ I Introduction:
Chalcogenide perovskites (CPs)[1, 2, 3, 4] have recently been emerged as a potential alternatives for the high-performing inorganic-organic halide perovskites (IOHPs)[5, 6, 7, 8]. While IOHPs face significant obstacles for their large-scale industrial applications owing to the toxicity of lead and instability due to the presence of organic cations[9, 10], CPs show promise because of their vast availability in the earth’s crust, non-toxic nature, and exceptional stability[1, 11, 12, 13]. In addition to these advantages, CPs exhibit favorable characteristics such as small electronic bandgap, high absorption coefficient, promising defect tolerance, good charge carrier mobility, and excellent power conversion efficiency (PCE)[2, 3, 4, 14, 15], which point toward their potential application in a variety of optoelectronic devices.
@@
Similar to the typical three-dimensional halide perovskites (HPs), the chemical formula of chalcogenide perovskites is expressed as ABX3[16, 4], where A and B stand for divalent alkali-earth metal cations (Ca2+, Sr2+, Ba2+) and tetravalent transition metal cations (Ti4+, Zr4+, Hf4+, and Sn4+), respectively, and X is typically a chalcogen anion such as S2- or Se2-. Numerous experimental and theoretical investigations have demonstrated the successful synthesis of chalcogenide perovskites and unveiled their intriguing properties[1, 11, 12, 13, 2, 3, 4, 14, 17, 18]. For example, Lelieveld et al.[17] had synthesized the distorted phase of CaZrS3, SrZrS3, BaZrS3, CaHfS3, SrHfS3, and BaHfS3 in 1980, while later, the needle-like phase[18, 19, 20] of SrZrS3, SrZrSe3, and SrHfSe3 CPs were also synthesized by various groups in the end of centuary and beginning of centuary. In 2016, Perera et al.[21] used high-temperature sulfurization of the oxides with CS2 to synthesize AZrS3 (A = Ba, Ca, and Sr). On the other hand, in 2015, Sun et al.[4] theoretically affirmed the formation of two distinct phases of CPs at room temperature: the needle-like phase (NH4CdCl3type) and the distorted phase (GdFeO3type), both sharing the orthorhombic structure with same space group (No. 62), and predicted them to be promising for application in solar cells. In addition, several first-principles DFT-based studies have been conducted, indicating that CPs exhibit interesting electronic and optical properties suitable for use in photovoltaic applications[14, 22, 16, 4, 15].
@@
Several studies have primarily focused on Zr- and Hf-based CPs[14, 22, 15], which demonstrate exciton binding energies ranging from 0.02 to 0.26 eV, comparable to or higher than those observed in conventional HPs (0.010.10 eV)[23, 24, 25]. However, the main drawback of Zr- and Hf-based CPs lies in their reduced charge carrier mobility (6.8477.59 cm2V-1s-1) and lower PCE (10.56%25.02%) compared to conventional HPs, which have charge carrier mobility and PCE in the range of 57290 cm2V-1s-1[26, 27] and 21.15%28.97%[28, 29], respectively. These limitations are majorly due to prominent polaronic effects, which hinder their effectiveness for solar cell applications. This makes exploration of alternatives for the B-site cation even more crucial in order to achieve high polaronic charge carrier mobility along with high PCE, thereby, enhancing the effectiveness of CPs for the solar cell applications.
@@
Recent first-principles calculations highlight the potential of Sn as B-site cation via. distorted SrSnX3 (X = S, Se) perovskites as promising photovoltaic materials solely based on their electronic and optical properties upto the HSE06 level[30, 31]. Experimental investigations have further substantiated this potential by successfully fabricating distorted CaSnS3 perovskite at 500°C[32]. Basera et al.[14] have furthermore provided valuable insights into the photovoltaic properties of distorted CaSnS3, as well as needle-like phases of BaSnS3 and SrSnS3 perovskites. However, the challenge posed by these needle-like CPs lies in their higher bandgap (1.912.04 eV) and lower theoretical efficiency (21.80%). While, in contrast, the bandgap of distorted CaSnS3 falls within the optimal range at 1.43 eV, with theoretical efficiency reaching up to 32.45%[14]. Motivated by these findings, we aim to conduct a comprehensive theoretical analysis of distorted Sn-based chalcogenides to elucidate their photovoltaic application suitable features.
@@
Henceforth, our endeavor has been to meticulously explore the optoelectronic, transport, excitonic, and polaronic properties of distorted Sn-based, ASnX3 (A = Ca, Sr, Ba; X = S, Se) chalcogenide perovskites utilizing a combination of density functional theory (DFT)[33, 34], density functional perturbation theory (DFPT)[35], and advanced many-body perturbation theory based methodologies like GW and BSE[36, 37]. Initially, the crystal structures are optimized using semilocal PBE[38] exchangecorrelation (xc) functional, and their stability has been examined. The electronic properties of the relaxed structures are calculated using HSE06[39] xc functional as well as G0W0@PBE[40, 41] methodology. The investigated CPs exhibit direct bandgaps ranging from 0.79 to 1.50 eV, which fall within the optimal range when compared to most lead-based HPs[23, 24, 28] as well as Zr- and Hf-based CPs[14, 22, 15]. Following this, optical properties and exciton binding energies () are determined through the Bethe-Salpeter equation (BSE)[42, 43] method. Furthermore, employing the DFPT technique, the ionic contribution to the dielectric function is computed, and the Fröhlich model is invoked to evaluate polaronic attributes such as carrier-phonon coupling strength and polaron mobility (21.33416.02 cm2V-1s-1). Conclusively, the estimation of the spectroscopic limited maximum efficiency (SLME)[44] between 24.20%31.20% utilizing the quasiparticle (QP) bandgap and absorption coefficient underscores ASnX3 as a material with significant potential for photovoltaic applications. This has been further confirmed by performing the conventional device (FTO/TiO2/ASnX3/Spiro-OMeTAD/Au) simulations using SCAPS-1D software [45, 46, 47, 48].
@@ II.1 Computational Details:
The phonon spectra were calculated using the DFPT method as implemented in the PHONOPY[53] package by considering a supercell. Since, GGA is known to underestimate the band gap, the electronic band structures were computed using the hybrid HSE06[39] xc functional as well as many-body perturbation theory (MBPT) based GW[40, 41] (G0W0@PBE) method. Note that the spin-orbit coupling (SOC) effect was not considered as it does not impact the bandgap (for details, see the Supplemental Material) of the considered systems. The effective mass was computed by SUMO[54] using a parabolic fitting of the band edges. We also carried out Bethe-Salpeter equation (BSE)[42, 43] based calculations on top of the single-shot GW(G0W0)@PBE to precisely estimate the optical properties, which takes explicitly into account the electron-hole interaction. Here, a -centered -grid and a converged 640 NBANDS were used for the GW-BSE calculations. The electron-hole kernel for the BSE calculations was generated by considering 24 occupied and 24 unoccupied bands. The VASPKIT[55] package was used to post-process the elastic and optical properties. The ionic contribution to the dielectric constant was also calculated using the DFPT method.
@@
where, and are the electronic (optical) and static (electronic + ionic) dielectric constants, and is the characteristic phonon angular frequency. The thermal "B" approach of Hellwarth et al.[58] is used to determine by taking the spectral average of the multiple phonon branches (for details, see the SM).
Within the framework of Fröhlich’s polaron model, the longitudinal optical phonons and the electron travelling through the lattice interact via the dimensionless Fröhlich parameter , which is expressed as[59],
where, is the permittivity of free space and is the carrier effective mass. One can also estimate the polaron energy () by knowing the value of using equation[22, 14]:
@@
where, is the charge of electron, , and and are the temperature-dependent variational parameters, and is a function of , , and (for details, see the SM).
@@ II.2 SCAPS-1D Numerical Simulations:
SCAPS-1D software is utilized for conducting numerical device simulations on solar cells to assess their performance and aid in design optimization. The tool was developed by Prof. M. Burgelman at the Department of Electronics and Information Systems (ELIS) of the University of Gent, Belgium[61, 62]. SCAPS solves the Poisson’s equation, which correlates the electrostatic potential to the overall charge density, along with continuity equations for electrons and holes in the conduction and valence bands, respectively. It can predict device characteristics such as current density-voltage curve, efficiency, energy bands, and other properties of the solar cell structure under illumination. Figure 5(a) shows the proposed perovskite solar cell (PSC) structure having the architecture of FTO/TiO2/Perovskite/Spiro-OMeTAD/Au. In this proposed structure, Spiro-OMeTAD iwa used as the hole transport layer (HTL), TiO2 was used as the electron transport layer (ETL), FTO (fluorine-doped tin oxide) as the transparent conducting oxide (TCO), and perovskite as the absorber layer. Gold (Au) having a work function of 5.1 eV was employed as the back metallic contact. The simulation parameters for the HTL, ETL and FTO, chosen based on theoretical and experimental results[45], are listed in Table S14 of the SM. Also, Table S15 shows the input parameters for the perovskite layer, estimated through our theoretical calculations (for details, see the SM). The simulations were conducted under the illumination of AM1.5G at 300K working temperature. The Poisson’s equation and the continuity equations of both the carriers (electrons and holes) for SCAPS-1D simulation are written as[46, 47],
Here is the electrostatic potential, is the permittivity of the material, and are electron and hole concentrations, and are donor and acceptor densities, and are electron and hole current densities, and are electron and hole distribution, is the recombination and is the optical generation rate.
@@
where, is the short circuit current density and is the reverse saturation current. The Fill Factor (FF) is given by[48]:
@@
where, is the maximum power of solar sell, and is the input solar power equivalent to the AM1.5G Sun spectrum.
@@ III Results and Discussions:
In this study, we undertook a systematic and thorough investigation into the distorted phases of Sn-based chalcogenide perovskites ASnX3 (A = Ca, Sr, Ba; X = S, Se) with the aim of exploring their potential optoelectronic features. The subsequent sections delve into the stability as well as the structural, electronic properties, transport phenomena, optical properties, excitonic dynamics, polaronic effects, and the spectroscopic limited maximum efficiency (SLME) of ASnX3 chalcogenide perovskites. Through this detailed examination and discussion, our aim is to establish a foundational understanding and provide insights to guide future experimental endeavors.
@@ III.1 Structural Properties:
The orthorhombic distorted crystal structure with the space group (No. 62)[17] of chalcogenide perovskites ASnX3 (A = Ca, Sr, Ba; X = S, Se) are presented in Figure 1(a) and Figure S1. The crystal structure of these compounds usually consists of four formula units, i.e., 20 atoms, of which 4 are Ca/Sr/Ba, 4 are Sn, and 12 are S or Se atoms. In this distorted phase, the alkaline earth elements (A-site cations) have 12-fold coordination, and they form cuboctahedrons with chalcogenides X (S or Se). On the other hand, the corner-sharing distorted octahedrons are produced due to the 6-fold coordination of Sn cation with X (S or Se) atoms in a distorted and tilted way. The lattice parameters and average bond lengths of the optimized crystal structures are calculated using the PBE xc functional and are tabulated in Table 1 and Table S1, respectively. It is found that the lattice parameters are in good agreement with previous theoretical and available experimental results[32, 30]. The crystallographic stability of these CPs is quantitatively identified by calculating the Goldschmidt tolerance factor (), the octahedral factor (), and the new tolerance factor ()[63, 64, 56] (for details, see the SM). From Table S2, it can be noticed that the , , and values of these investigated CPs are in the range of 0.8790.964, 0.3480.375, and 4.1744.721, respectively, confirming the reasonable structural stability of these distorted CPs[63, 16]. Furthermore, the decomposition energy is also calculated for these materials using the PBE xc functional to ascertain their thermodynamic stability (for details, see Table S3 of the SM). The negative decomposition energy reveals these perovskites to be unstable at 0 K, but, there remains a possibility for them to stabilize at higher temperatures, as also demonstrated by Hamza et al. by synthesizing CaSnS3 perovskite at 500° C[32], which is an indicator of their higher temperature stability. Further, the dynamical and mechanical stability of the studied CPs are examined.
@@
In Figure 1(b)-(c), the phonon spectra of SrSnS3 and BaSnS3 CPs have been depicted, which validates their dynamical stability at 0 K. The rest of the compounds are found to be unstable at 0 K but they may become dynamically stable at higher temperatures, as discussed before. To investigate the mechanical stability, the second-order elastic coefficients () of these CPs are calculated using the energy-strain approach[65] (for details, see the SM). The computed values of these compounds are listed in Table S4, and they are found to satisfy the Born stability criteria[65]. Using these elastic coefficients, the bulk modulus (), shear modulus (), Young’s modulus (), and Poisson’s ratio () of these materials[66, 67] are investigated (for details, see the SM). The fragility of the materials is studied in terms of Pugh’s suggested ratio ()[68] and Poisson’s ratio (). The calculated values of (> 1.75) and (> 0.26) reveal that the examined CPs are ductile in nature (see Table S4). Additionally, the longitudinal (), transverse (), and average () elastic wave velocities and the Debye temperature ()[69] are calculated (see Table S5), as they hold crucial significance for flexible optoelectronic applications.

### v2

# Post-Transition Metal Sn-Based Chalcogenide Perovskites: A Promising Lead-Free and Transition Metal Alternative for Stable, High-Performance Photovoltaics

@@ Abstract
Chalcogenide perovskites (CPs) have emerged as promising materials for optoelectronic applications due to their stability, non-toxicity, small bandgaps, high absorption coefficients, and defect tolerance. Although transition metal-based CPs, particularly those incorporating Zr and Hf, have been well-studied, they often exhibit higher bandgaps, lower charge carrier mobility, and reduced efficiency compared to lead-based halide perovskites (HPs). Tin (Sn), a post-transition metal with a similar oxidation state (+4) as Zr and Hf in ABX3 structures but with different valence characteristics, remains underexplored in CPs. Given the influence of valence states on material properties, Sn-based CPs are of great interest. This study employs density functional theory (DFT), density functional perturbation theory (DFPT), and many-body perturbation theory (GW and BSE) to investigate a series of distorted Sn-based CPs (ASnX3, A = Ca, Sr, Ba; X = S, Se). Our results demonstrate that these perovskites are mechanically stable and exhibit lower direct G0W0 bandgaps (0.79–1.50 eV) compared to their Zr- and Hf-based counterparts. Analysis of carrier-phonon interactions reveals that the charge-separated polaronic state is less stable than the bound exciton state in these materials. Additionally, polaron-assisted charge carrier mobilities for electrons (21.33–416.02 cm2V-1s-1) and holes (7.02–260.69 cm2V-1s-1) are comparable to or higher than those in lead-based HPs and significantly exceed those of Zr- and Hf-based CPs, owing to reduced carrier-phonon coupling. The estimated spectroscopic limited maximum efficiency (24.2%–31.2%)—confirmed through perovskite solar cell (PSC) simulations using SCAPS-1D software—indicates that these materials are promising candidates for photovoltaic applications. Overall, this study highlights the potential of Sn as a superior alternative to transition metals in CPs, particularly for photovoltaic applications, where smaller electronegativity differences lead to reduced bandgaps and reduced polaronic effects improved charge carrier mobility. These findings are expected to stimulate further experimental investigation into Sn-based CPs.
@@ I Introduction:
Chalcogenide perovskites (CPs)[1, 2, 3, 4] have recently emerged as a potential alternative for the high-performing inorganic-organic halide perovskites (IOHPs)[5, 6, 7, 8]. While IOHPs face significant obstacles for their large-scale industrial applications owing to the toxicity of lead and instability due to the presence of organic cations[9, 10], CPs show promise because of their vast availability in the earth’s crust, non-toxic nature, and exceptional stability[1, 11, 12, 13]. In addition to these advantages, CPs exhibit favorable characteristics such as small electronic bandgap, high absorption coefficient, promising defect tolerance, good charge carrier mobility, and excellent power conversion efficiency (PCE)[2, 3, 4, 14, 15], which point toward their potential application in a variety of optoelectronic devices.
@@
Similar to the typical three-dimensional halide perovskites (HPs), the chemical formula of chalcogenide perovskites is expressed as ABX3[16, 4], where A and B stand for divalent alkali-earth metal cations (Ca2+, Sr2+, Ba2+) and tetravalent transition metal cations (Ti4+, Zr4+, and Hf4+), respectively, and X is typically a chalcogen anion such as S2- or Se2-. Numerous experimental and theoretical investigations have demonstrated the successful synthesis of chalcogenide perovskites and unveiled their intriguing properties[1, 11, 12, 13, 2, 3, 4, 14, 17, 18]. For example, Lelieveld et al.[17] had synthesized the distorted phase of CaZrS3, SrZrS3, BaZrS3, CaHfS3, SrHfS3, and BaHfS3 in 1980, while later, the needle-like phase[18, 19, 20] of SrZrS3, SrZrSe3, and SrHfSe3 CPs were also synthesized by various groups in the end of century and beginning of century. In 2016, Perera et al.[21] used high-temperature sulfurization of the oxides with CS2 to synthesize AZrS3 (A = Ba, Ca, and Sr). On the other hand, in 2015, Sun et al.[4] theoretically affirmed the formation of two distinct phases of CPs at room temperature: the needle-like phase (NH4CdCl3type) and the distorted phase (GdFeO3type), both sharing the orthorhombic structure with same space group (No. 62), and predicted them to be promising for application in solar cells. In addition, several first-principles DFT-based studies have been conducted, indicating that CPs exhibit interesting electronic and optical properties suitable for use in photovoltaic applications[14, 22, 16, 4, 15].
@@
Several studies have primarily focused on Zr ([Kr]4d25s2) and Hf ([Xe]4f145d26s2) transition metal-based CPs, where the valence band configurations involve - and - orbitals[14, 22, 15]. These materials are demonstrated to exhibit exciton binding energies ranging from 0.02 to 0.26 eV, which are comparable to or higher than those observed in conventional lead-based HPs (0.010.10 eV)[23, 24, 25]. However, the main drawback of these CPs lies in their reduced charge carrier mobility (6.8477.59 cm2V-1s-1) and lower PCE (10.56%25.02%) compared to conventional HPs, which have charge carrier mobility and PCE in the range of 57290 cm2V-1s-1[26, 27] and 21.15%28.97%[28, 29], respectively. These limitations are majorly due to prominent polaronic effects and relatively higher bandgaps, which hinder their effectiveness for solar cell applications. This makes exploration of alternatives for the B-site cation even more crucial in order to achieve high polaronic charge carrier mobility along with high PCE, thereby enhancing the effectiveness of CPs for solar cell applications.
@@
The polaronic effects are more pronounced in transition metal-based CPs. However, if the post-transition metal Sn ([Kr]4d105s25p2), which also exhibits a oxidation state, is used in place of transition metals like Zr and Hf, the polaronic effects could be mitigated due to different valence states. Additionally, a lower difference in the electronegativity of Sn and X (S, Se) may offer the potential for achieving appropriate bandgaps for photovoltaic applications. Recent first-principles calculations highlight the potential of Sn as a B-site cation via. distorted SrSnX3 (X = S, Se) perovskites as promising photovoltaic materials solely based on their electronic and optical properties up to the HSE06 level[30, 31]. Experimental investigations have substantiated this potential by fabricating distorted CaSnS3 perovskite at 500°C[32]. Basera et al.[14] have furthermore provided valuable insights into the photovoltaic properties of distorted CaSnS3, as well as needle-like phases of BaSnS3 and SrSnS3 perovskites. However, the challenge posed by these needle-like CPs lies in their higher bandgap (1.912.04 eV) and lower theoretical efficiency (21.80%). In contrast, the bandgap of distorted CaSnS3 falls within the optimal range at 1.43 eV, with theoretical efficiency reaching up to 32.45%[14]. Motivated by these findings, we aim to conduct a comprehensive theoretical analysis of post-transition metal Sn-based distorted CPs to elucidate their suitable photovoltaic application features.
@@
Henceforth, we have endeavored to meticulously explore the optoelectronic, transport, excitonic, and polaronic properties of distorted Sn-based, ASnX3 (A = Ca, Sr, Ba; X = S, Se) CPs utilizing a combination of density functional theory (DFT)[33, 34], density functional perturbation theory (DFPT)[35], and advanced many-body perturbation theory based methodologies like GW and BSE[36, 37]. Initially, the crystal structures are optimized using semilocal PBE[38] exchangecorrelation (xc) functional, and their stability has been examined. The electronic properties of the relaxed structures are calculated using HSE06[39] xc functional as well as G0W0@PBE[40, 41] methodology. The investigated CPs exhibit direct bandgaps ranging from 0.79 to 1.50 eV, which fall within the similar range to that of conventional lead-based HPs[23, 28] and are lower than the bandgaps observed in Zr- and Hf-based CPs[14, 22, 15]. Following this, optical properties and exciton binding energies () are determined through the Bethe-Salpeter equation (BSE)[42, 43] method. Furthermore, employing the DFPT technique, the ionic contribution to the dielectric function is computed, and the Fröhlich model is invoked to evaluate polaronic attributes such as carrier-phonon coupling strength and polaron mobility (21.33416.02 cm2V-1s-1). Conclusively, the estimation of the spectroscopic limited maximum efficiency (SLME)[44] between 24.20%31.20% utilizing the quasiparticle (QP) bandgap and absorption coefficient underscores ASnX3 as a material with significant potential for photovoltaic applications. This has been further confirmed by performing the conventional device (FTO/TiO2/ASnX3/Spiro-OMeTAD/Au) simulations using SCAPS-1D software [45, 46, 47, 48].
@@ II.1 Computational Details:
The phonon spectra were calculated using the DFPT method as implemented in the PHONOPY[53] package by considering a supercell. Since, GGA is known to underestimate the bandgap, the electronic band structures were computed using the hybrid HSE06[39] xc functional as well as many-body perturbation theory (MBPT) based GW[40, 41] (G0W0@PBE) method. Note that the spin-orbit coupling (SOC) effect was not considered as it does not impact the bandgap (for details, see the Supplemental Material) of the considered systems. The effective mass was computed by SUMO[54] using a parabolic fitting of the band edges. We also carried out Bethe-Salpeter equation (BSE)[42, 43] based calculations on top of the single-shot GW(G0W0)@PBE to precisely estimate the optical properties, which takes explicitly into account the electron-hole interaction. Here, a -centered -grid and a converged 640 NBANDS were used for the GW-BSE calculations. The polarizability calculations were performed using a frequency grid of 60 points to ensure accurate results. In our GW calculations, we used the plasmon-pole approximation (PPA) to model the frequency dependence of the dielectric function (ALGO = GW0). Additionally, the plane-wave cutoff for the response function was set to ENCUTGW = 267 eV (2/3 of ENCUT) for both GW and BSE calculations. The electron-hole kernel for the BSE calculations was generated by considering 24 occupied and 24 unoccupied bands. The VASPKIT[55] package was used to post-process the elastic and optical properties. The ionic contribution to the dielectric constant was also calculated using the DFPT method.
@@
where, and are the electronic (optical) and static (electronic + ionic) dielectric constants, and is the characteristic phonon angular frequency. The thermal "B" approach of Hellwarth et al.[58] is used to determine by taking the spectral average of the multiple phonon branches (for details, see the Supplemental Material).
Within the framework of Fröhlich’s polaron model, the longitudinal optical phonons and the electron traveling through the lattice interact via the dimensionless Fröhlich parameter , which is expressed as[59],
where is the permittivity of free space and is the carrier effective mass. One can also estimate the polaron energy () by knowing the value of using equation[22, 14]:
@@
where is the charge of the electron, , and and are the temperature-dependent variational parameters, and is a function of , , and (for details, see the Supplemental Material).
@@ II.2 SCAPS-1D Numerical Simulations:
SCAPS-1D software is utilized to conduct numerical device simulations on solar cells to assess their performance and aid in design optimization. The tool was developed by Prof. M. Burgelman at the Department of Electronics and Information Systems (ELIS) of the University of Gent, Belgium[61, 62]. SCAPS solves Poisson’s equation, which correlates the electrostatic potential to the overall charge density, along with continuity equations for electrons and holes in the conduction and valence bands, respectively. It can predict device characteristics such as current density-voltage curve, efficiency, energy bands, and other properties of the solar cell structure under illumination. Figure 5(a) shows the proposed perovskite solar cell (PSC) structure having the architecture of FTO/TiO2/Perovskite/Spiro-OMeTAD/Au. In this proposed structure, Spiro-OMeTAD was used as the hole transport layer (HTL), TiO2 was used as the electron transport layer (ETL), FTO (fluorine-doped tin oxide) as the transparent conducting oxide (TCO), and perovskite as the absorber layer. Gold (Au), having a work function of 5.1 eV, was employed as the back metallic contact. The simulation parameters for the HTL, ETL, and FTO, chosen based on theoretical and experimental results[45], are listed in Table S15 of the Supplemental Material. Table S16 also shows the input parameters for the perovskite layer, estimated through our theoretical calculations (for details, see the Supplemental Material). The simulations were conducted under the illumination of AM1.5G at 300K working temperature. The Poisson’s equation and the continuity equations of both the carriers (electrons and holes) for SCAPS-1D simulation are written as[46, 47],
Here is the electrostatic potential, is the permittivity of the material, and are electron and hole concentrations, and are donor and acceptor densities, and are electron and hole current densities, and are electron and hole distribution, is the recombination, and is the optical generation rate.
@@
where is the short circuit current density and is the reverse saturation current. The Fill Factor (FF) is given by[48]:
@@
where is the maximum power of the solar cell, and is the input solar power equivalent to the AM1.5G Sun spectrum.
@@ III Results and Discussions:
In this study, we undertook a systematic and thorough investigation into the distorted phases of post-transition metal Sn-based chalcogenide perovskites ASnX3 (A = Ca, Sr, Ba; X = S, Se) to explore their potential optoelectronic features. The subsequent sections delve into the stability as well as the structural, electronic properties, transport phenomena, optical properties, excitonic dynamics, polaronic effects, and the spectroscopic limited maximum efficiency (SLME) of ASnX3 chalcogenide perovskites. This detailed examination and discussion aims to establish a foundational understanding and provide insights to guide future experimental endeavors.
@@ III.1 Structural Properties:
The orthorhombic distorted crystal structure with the space group (No. 62)[17] of chalcogenide perovskites ASnX3 (A = Ca, Sr, Ba; X = S, Se) are presented in Figure 1(a) and Figure S1. The crystal structure of these compounds usually consists of four formula units, i.e., 20 atoms, of which 4 are Ca/Sr/Ba, 4 are Sn, and 12 are S or Se atoms. In this distorted phase, the alkaline earth elements (A-site cations) have 12-fold coordination, and they form cuboctahedrons with chalcogenides X (S or Se). On the other hand, the corner-sharing distorted octahedrons are produced due to the 6-fold coordination of Sn cation with X (S or Se) atoms in a distorted and tilted way. The lattice parameters of the optimized crystal structures are calculated using the PBE xc functional and are tabulated in Table 1. It is found that the lattice parameters are in good agreement with previous theoretical and available experimental results[32, 30]. However, for CaSnS3, there is a discrepancy in the c-lattice constant between the theoretical and experimental values, which could influence the bandgap, as variations in bond lengths and angles modify the electronic band dispersion. In addition, Table S1 presents detailed information on the octahedral distortions of SnX6 octahedra in ASnX3 (A = Ca, Sr, Ba; X = S, Se) CPs, including parameters such as average bond length, bond angle variance, polyhedral volume, and distortion index. The results indicate that the SnX6 octahedra in BaSnX3 (X = S, Se) undergo greater lattice distortion than those in other compounds, likely due to the larger ionic radius of Ba2+, which influences the local bonding environment around the octahedra.
The crystallographic stability of these CPs is quantitatively identified by calculating the Goldschmidt tolerance factor (), the octahedral factor (), and the new tolerance factor ()[63, 64, 56] (for details, see the Supplemental Material). From Table S2, it can be noticed that the , , and values of these investigated CPs are in the range of 0.8790.964, 0.3480.375, and 4.1744.721, respectively, confirming the reasonable structural stability of these distorted CPs[63, 16]. Furthermore, the decomposition energy is also calculated for these materials using the PBE xc functional to ascertain their thermodynamic stability (for details, see Table S3 of the Supplemental Material). The negative decomposition energy reveals these perovskites to be unstable at 0 K. Still, there remains a possibility for them to stabilize at higher temperatures, as also demonstrated by Hamza et al. by synthesizing CaSnS3 perovskite at 500° C[32], which is an indicator of their higher temperature stability. Further, the dynamical and mechanical stability of the studied CPs are examined.
@@
In Figure 1(b)-(c), the phonon spectra of SrSnS3 and BaSnS3 CPs have been depicted, which validates their dynamical stability at 0 K. The rest of the compounds are found to be unstable at 0 K, but they may become dynamically stable at higher temperatures, as discussed before. To investigate the mechanical stability, the second-order elastic coefficients () of these CPs are calculated using the energy-strain approach[65] (for details, see the Supplemental Material). The computed values of these compounds are listed in Table S4, and they are found to satisfy the Born stability criteria[65]. Using these elastic coefficients, the bulk modulus (), shear modulus (), Young’s modulus (), and Poisson’s ratio () of these materials[66, 67] are investigated (for details, see the Supplemental Material). The fragility of the materials is studied in terms of Pugh’s suggested ratio ()[68] and Poisson’s ratio (). The calculated values of (> 1.75) and (> 0.26) reveal that the examined CPs are ductile in nature (see Table S4). Additionally, the longitudinal (), transverse (), and average () elastic wave velocities and the Debye temperature ()[69] are calculated (see Table S5), as they hold crucial significance for flexible optoelectronic applications.